A hollow copper wire with an inner diameter of 0.60 mm and an outer diameter of 2.1 mm carries a current of 10 A .

What is the current density in the wire?
Answer #1

The cross-sectional area of the wire is, A= (- ) =-|(2.1x109 m)?- (0.60x10^m)*] =1.27x10- m? The current density current per
Answer #2

A= π(r2-r1)

r2= 2.1/2 = 1.15

r1= 0.6/2 = 0.3

A= π((1.05 * 10-3)2-(0.3*10-3)2

A= 3.18 * 10-6 m2


J = I/A

J = 10/(3.18 * 10-6)

J = 3.14 * 106 A/m2

source: I did a very similar question.
answered by: Pill Dickle